OpenStudy (anonymous):

The force F = 4 i + 2 j moves an object 3 ft along the x-axis in the positive direction. Find the work done if the unit of force is the pound.

OpenStudy (xishem):

Work is defined as:\[W=\vec{F}\cdot \vec{d}\]

OpenStudy (akashdeepdeb):

Work done is scalar = Force . Displacement = | Force | * |Displacement| * cos theta Theta = angle between force and displacement. Or do it by the dot product method. [Which is much faster] If Force [F] = Fx + Fy + Fz Displacement [S] = Sx + Sy + Sz Then work done is their dot product which is = Fx*Sx + Fy*Sy + Fz*Sz [That is multiply their magnitudes] Got it? :D

OpenStudy (akashdeepdeb):

Here you have F = 4i + 2j OR 4x + 2y Displacement is along x axis so S = 3x or 3i Now simply find the dot product by the latter method! :D

OpenStudy (anonymous):

12 right for the dot method?

OpenStudy (akashdeepdeb):

correct